OpenStudy (alesia/bailey):

Using the given information, write an equation in slope-intercept form. then graph. (-1/2,3/4)(-3,2)

OpenStudy (haleyelizabeth2017):

First, you want to find the slope. Do you know how to do that?

OpenStudy (zpupster):

y = mx + b m= slope and b is y intercept m = y2-y1/x2-x1

OpenStudy (haleyelizabeth2017):

You will want to find the slope before we can do anything else :)

OpenStudy (alesia/bailey):

okay.. don't you stack and subtract???

OpenStudy (haleyelizabeth2017):

so, you use the formula @zpupster gave you. m=y2-y1/x2-x1. (x1,y1), (x2,y2) and you plug in the points.

OpenStudy (haleyelizabeth2017):

So it will look like (2-3/4)/(-3-(-1/2)) and you need to solve. Remember, subtracting a negative is the same as adding a positive. :)

OpenStudy (alesia/bailey):

okay hang on :P

OpenStudy (haleyelizabeth2017):

okay :)

OpenStudy (haleyelizabeth2017):

Did you get it?

OpenStudy (haleyelizabeth2017):

It should turn out as a decimal.....

OpenStudy (alesia/bailey):

how do u subtract the whole number and fractions?? I suck with fractions

OpenStudy (haleyelizabeth2017):

so turn them into decimals. so 2-0.75/-3+0.5

OpenStudy (haleyelizabeth2017):

1.25/-2.5=-0.5 :) now plug that into the equation as m :) Do that and then tell me what you get for the equation so far.

OpenStudy (alesia/bailey):

wait okay, so subtracting I got 1.25/3.5

OpenStudy (haleyelizabeth2017):

-2.5, not 3.5 :) remember, the 3 was negative

OpenStudy (alesia/bailey):

oh okay. xD srry, im just now getting out of school o3o

OpenStudy (haleyelizabeth2017):

Here, I will just show you the steps..... y=-0.5x+b now we will plug in one of the points to find b. (2)=-0.5(-3)+b 2=1.5+b -1.5 0.5=b now plug b back into the equation y=-0.5x+0.5 So I hope you can tell what I did :) Good luck!